# Near-Optimal Regret for Efficient Stochastic Combinatorial Semi-Bandits

### v1

@@ Appendix A Techinical Proofs of Section 3
$\leq\sum_{i\notin A^{*}}\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbf{E}\left[\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{2^{\ell+1}\Delta_{i,\ell,n}}{k_{i,\ell,n}}\;\middle|\;i\in A_{t}\right]$
$=2^{\ell+1}\sum_{i\notin A^{*}}\mathbf{E}\left[\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\;\middle|\;i\in A_{t}\right],$ (5)
where the first inequality is from Lemma 3. To avoid redundant computations for regret at eq. 5, we could arrange the sequence of actions that maximizes the value. Moreover, applying Lemma 9, we have the following lemma:
@@
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\right]\Delta_{A_{t}}\right]$
@@
$\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}},$
@@
$\phantom{{}={}}\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}$
@@
where eq. 6 is due to $\mathbb{1}\left[x\leq y<z\right]=\mathbb{1}\left[y\geq x\right]-\mathbb{1}\left[y\geq z\right]$ for any $y$ if $x<z$.
@@
Combining this with eq. 5 and eq. 6, we have
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\land\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$
@@
Perceiving $\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}$ as the variable $x$, we could bound the first term by the integral of $f(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x^{2}}$ over $x>\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}>\sqrt{\frac{em}{kT}}=\sqrt{e\delta}$, while the second term is bounded by the maximum of $g(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x}$ over $x>\sqrt{\frac{em}{kT}}$. $f(x)$ has been plotted in Figure 2. It is easy to verify that $f(x)$ is monotonically decreasing and positive when $x>\sqrt{\frac{e\delta}{k^{2}}}=\sqrt{\frac{em}{k^{3}T}}$. Additionally we have $\int f(x)dx=\frac{\ln\frac{\delta}{e^{2}k^{2}}-2\ln x}{x}$. Thus the summation could be bounded by the integral of $f(x)$ on $(\sqrt{\frac{em}{kT}},\infty)$. Also $g(x)$ is monotonically decreasing when $x>\sqrt{\frac{em}{kT}}$, which gives an upper bound for the second term $\frac{k_{i,\ell,1}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,1}^{2}}{k_{i,\ell,1}^{2}}\right)}{\Delta_{i,\ell,1}}$. Thus we have
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\right]\Delta_{A_{t}}\right]$ $\leq 1540\sum_{i\notin A^{*}}\left(\int_{\sqrt{\frac{em}{kT}}}^{\infty}f(x)dx+\frac{\ln ek^{2}}{\sqrt{\frac{em}{kT}}}\right)$
$=1540\sum_{i\notin A^{*}}\left(\frac{\ln\frac{\delta}{e^{2}k^{2}}-2\ln x}{x}\Big{|}_{\sqrt{\frac{em}{kT}}}^{\infty}+\ln ek^{2}\sqrt{\frac{kT}{em}}\right)$
$=1540\sum_{i\notin A^{*}}\left(\ln e^{4}k^{4}\sqrt{\frac{kT}{em}}\right)=O\left(\log k\sqrt{kmT}\right).$
@@
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\left[\Delta_{A_{t}}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},\sqrt{\frac{ekm}{T}}\right\}\right]\Delta_{A_{t}}\right]$
$=\sum_{\ell=0}^{\lfloor\log k\rfloor}\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\land\Delta_{A_{t}}>\max\left\{8\sum_{j\in A^{*}}\zeta_{j},\sqrt{\frac{ekm}{T}}\right\}\right]\Delta_{A_{t}}\right]$
$=O\left((\log k)^{2}\sqrt{kmT}\right).$
Bring this back to eq. 1, we could obtain $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\right]=O\left((\log k)^{2}\sqrt{kmT}\right)$.
@@
$=O\left(\frac{(\log k)^{2}}{p^{*}}\sqrt{kmT}\right),$ (7)
where eq. 7 applies Theorem 1 to this case.
For the conjunctive case, $r(A_{t},\bm{\mu})=\prod_{i\in A_{t}}\mu_{i}$. The rest analysis follows the same pattern as the conjunctive case.
@@ B.1 The CUCB algorithm
In Algorithm 3, if all $\omega_{i}(t)$ are less than $\left(\frac{1}{k}-\frac{\gamma}{m}\right)\sum_{i=1}^{m}\frac{\omega_{i}(t)}{1-\gamma}$, then every $p_{i}(t)$ calculated at Line 9 is less than $1$ and $S_{0}(t)=\emptyset$. Otherwise it chooses a appropriate threshold $\alpha_{t}$ that the temporal weight $\omega_{i}^{\prime}(t)$ is set to $\alpha_{t}$ for $i\in S_{0}(t)$ and $\omega_{i}(t)$ for $i\notin S_{0}(t)$. Thus $p_{i}(t)=1$ for $i\in S_{0}(t)$, which means that $S_{0}(t)\subseteq S(t)$. And $\omega_{i}(t)$ is not updated for $i\in S_{0}(t)$ since it is large to some extent.
@@ C.1 Comparison with baselines and beyond
As described in the main text, in the synthetic setting, we simulate a combinatorial bandit problem with $m=30$ base arms and cardinality $k=10$. The initial value for each arm is independently drawn from uniform distributions $\mathtt{U}(0,0.1)$ and $\mathtt{U}(0.5,0.6)$. Each algorithm is run for $100000$ rounds, and all results are averaged over $10$ independent runs.
@@
To match the synthetic regimes used in our experiments, the raw affinity scores are min-max normalized and linearly rescaled into two ranges: a low-probability regime $[0,0.1]$ and a high-probability regime $[0.5,0.6]$. Specifically, let $\min$ and $\max$ denote the minimum and maximum dot products across all user–item pairs. The scores are transformed as follows:
@@
$\text{score}_{\text{high}}(u,a)$ $=0.5+0.1\cdot\frac{\mathbf{u}^{\top}\mathbf{a}-\min}{\max-\min}.$ (10)
The resulting values are interpreted as the initial click probabilities for base arms. For each regime, we independently sample $30$ such values as base arms using a fixed random seed ($2025$) to ensure reproducibility.
@@
Metric Algorithm Range $[0,0.1]$ Range $[0.5,0.6]$

### v2

@@ Appendix A Techinical Proofs of Section 3
$=2^{\ell+1}\sum_{i\notin A^{*}}\mathbf{E}\left[\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[A_{t}=A_{i,\ell,n}\right]\cdot\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\;\middle|\;i\in A_{t}\right],$ (5)
where the first inequality is from Lemma 3. To avoid redundant computations for regret at Equation 5, we could arrange the sequence of actions that maximizes the value. Moreover, applying Lemma 9, we have the following lemma:
@@
$\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[A_{t}=A_{i,\ell,n}\right]\cdot\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}},$
@@
$\phantom{{}={}}\sum_{t=1}^{T}\sum_{n=1}^{N_{i,\ell}}\mathbb{1}\left[A_{t}=A_{i,\ell,n}\right]\cdot\mathbb{1}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{(\sqrt{2})^{\ell}}{16}\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}\right]\cdot\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}$
@@
where Equation 6 is due to $\mathbb{1}\left[x\leq y<z\right]=\mathbb{1}\left[y\geq x\right]-\mathbb{1}\left[y\geq z\right]$ for any $y$ if $x<z$.
@@
Combining this with Equation 5 and Equation 6, we have
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$
@@
Perceiving $\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}$ as the variable $x$, we could bound the first term by the integral of $f(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x^{2}}$ over $x>\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}>(\log k)\sqrt{\frac{em}{kT}}$, while the second term is bounded by the maximum of $g(x)=\frac{\ln\left(\frac{k^{2}x^{2}}{\delta}\right)}{x}$ over $x>(\log k)\sqrt{\frac{em}{kT}}$. $f(x)$ has been plotted in Figure 2. It is easy to verify that $f(x)$ is monotonically decreasing and positive when $x>\sqrt{\frac{e\delta}{k^{2}}}=(\log k)\sqrt{\frac{em}{k^{3}T}}$. Additionally we have $\int f(x)dx=\frac{\ln\frac{\delta}{e^{2}k^{2}}-2\ln x}{x}$. Thus the summation could be bounded by the integral of $f(x)$ on $((\log k)\sqrt{\frac{em}{kT}},\infty)$. Also $g(x)$ is monotonically decreasing when $x>(\log k)\sqrt{\frac{em}{kT}}$, which gives an upper bound for the second term $\frac{k_{i,\ell,1}\ln\left(\frac{k^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,1}^{2}}{k_{i,\ell,1}^{2}}\right)}{\Delta_{i,\ell,1}}$. Thus we have
$\sum_{t=1}^{T}\mathbf{E}\left[\mathbb{1}\left[A_{t}\in\mathcal{A}_{\ell}\right]\Delta_{A_{t}}\right]$ $\leq 1540\sum_{i\notin A^{*}}\left(\int_{(\log k)\sqrt{\frac{em}{kT}}}^{\infty}f(x)dx+\frac{\ln ek^{2}}{(\log k)\sqrt{\frac{em}{kT}}}\right)$
$=1540\sum_{i\notin A^{*}}\left(\frac{\ln\frac{\delta}{e^{2}k^{2}}-2\ln x}{x}\Big|_{(\log k)\sqrt{\frac{em}{kT}}}^{\infty}+\frac{\ln ek^{2}}{\log k}\sqrt{\frac{kT}{em}}\right)$
$=1540\sum_{i\notin A^{*}}\left(\frac{\ln e^{4}k^{4}}{\log k}\sqrt{\frac{kT}{em}}\right)=O\left(\sqrt{kmT}\right).$
@@
$\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$ $=\sum_{\ell=0}^{\lfloor\log k\rfloor}\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\Delta_{A_{t}}\right]$
$=O\left((\log k)\sqrt{kmT}\right).$
Bring this back to Equation 1, we could obtain $\mathbf{E}\left[R(T)\right]=\mathbf{E}\left[\sum_{t=1}^{T}\Delta_{A_{t}}\right]=O\left((\log k)\sqrt{kmT}\right)$.
The case where $k>\frac{m}{2}$
In this case, we need to prove that the regret of CMOSS is $O\left((m-k)\sqrt{\log k\log(m-k)T}\right)$. Here $\delta$ is optimized to be $\frac{(m-k)^{2}\log k\log(m-k)}{k^{2}T}$. Now we let $\mathcal{A}_{B}=\left\{A\in\mathcal{A}\;\middle|\;\Delta_{A}>\max\left\{8\sum_{j=1}^{k}\zeta_{j},(m-k)\sqrt{\frac{e\log k\log(m-k)}{T}}\right\}\right\}$. And we decompose the overall regret as follows:
$\sum_{t=1}^{T}\mathbf{E}\left[\Delta_{A_{t}}\right]$ $=\sum_{t=1}^{T}\mathbf{E}\left[\left[\Delta_{A_{t}}\leq(m-k)\sqrt{\frac{e\log k\log(m-k)}{T}}\right]\Delta_{A_{t}}\right]+\sum_{t=1}^{T}\mathbf{E}\left[\left[\Delta_{A_{t}}\leq 8\sum_{j=1}^{k}\zeta_{j}\right]\Delta_{A_{t}}\right]$
$+\sum_{t=1}^{T}\mathbf{E}\left[\left[\Delta_{A_{t}}>\max\left\{8\sum_{j=1}^{k}\zeta_{j},(m-k)\sqrt{\frac{e\log k\log(m-k)}{T}}\right\}\right]\Delta_{A_{t}}\right]$
$\leq(m-k)\sqrt{e\log k\log(m-k)T}+8\sum_{j=1}^{k}\mathbf{E}\left[\zeta_{j}\right]\cdot T+\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{B}\right]\Delta_{A_{t}}\right].$ (7)
From Lemma 8, the second term could be bounded: $8\sum_{j=1}^{k}\mathbf{E}\left[\zeta_{j}\right]\cdot T\leq 8k\sqrt{\delta}\cdot T=8(m-k)\sqrt{\log k\log(m-k)T}$.
Since now the range of $k_{A}$ changes to $1\leq k_{A}\leq m-k$, from Lemma 2, we could divide the action set $\mathcal{A}$ into $1+\lfloor\log(m-k)\rfloor$ exhaustive and mutually exclusive sets $\left(\mathcal{A}_{\ell}\right)_{0\leq\ell\leq\lfloor\log(m-k)\rfloor}$, where
Furthermore, we could similarly prove Lemma 3 in this case: for every $0\leq\ell\leq\lfloor\log(m-k)\rfloor$,
$\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\leq\frac{2^{\ell+1}}{k_{A_{t}}}\sum_{i\notin A^{*}}\left[\hat{\mu}_{i,t}+\rho_{i,t}\geq\mu_{i}+\frac{\left(\sqrt{2}\right)^{\ell}}{16}\cdot\frac{\Delta_{A_{t}}}{k_{A_{t}}}\land i\in A_{t}\right].$ (8)
Next we give the actions in every $\mathcal{A}_{\ell}$ a regret upper bound. That is, for every $0\leq\ell\leq\lfloor\log(m-k)\rfloor$, we need to prove
$\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]=O\left((m-k)\sqrt{\frac{T\log k}{\log(m-k)}}\right).$
Following the same process of the proof of Lemma 4, by slightly changing the upper bound of $k_{A}$ from $k$ to mkm-k, we have that, for every $0\leq\ell\leq\lfloor\log(m-k)\rfloor$,
$\phantom{{}={}}\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\right]\Delta_{A_{t}}\right]$
$\leq 1540\sum_{i\notin A^{*}}\left(\sum_{n=2}^{N_{i,\ell}}\frac{k_{i,\ell,n}^{2}\ln\left(\frac{(m-k)^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,n}^{2}}{k_{i,\ell,n}^{2}}\right)}{\Delta_{i,\ell,n}^{2}}\left(\frac{\Delta_{i,\ell,n}}{k_{i,\ell,n}}-\frac{\Delta_{i,\ell,n-1}}{k_{i,\ell,n-1}}\right)+\frac{k_{i,\ell,1}\ln\left(\frac{(m-k)^{2}}{\delta}\cdot\frac{\Delta_{i,\ell,1}^{2}}{k_{i,\ell,1}^{2}}\right)}{\Delta_{i,\ell,1}}\right).$
Next we consider the properties of the function $f(x)=\frac{\ln\left(\frac{(m-k)^{2}x^{2}}{\delta}\right)}{x^{2}},x>\frac{\Delta_{i,\ell,1}}{k_{i,\ell,1}}>\sqrt{\frac{e\log k\log(m-k)}{T}}=\frac{k\sqrt{e\delta}}{m-k}$. It is easy to verify that $f(x)$ is monotonically decreasing and positive when $x>\sqrt{\frac{e\delta}{(m-k)^{2}}}=\frac{\sqrt{e\delta}}{m-k}$. Additionally we have $\int f(x)dx=\frac{\ln\frac{\delta}{e^{2}(m-k)^{2}}-2\ln x}{x}$. Also $g(x)=\frac{\ln\left(\frac{(m-k)^{2}x^{2}}{\delta}\right)}{x}$ is monotonically decreasing when $x>\frac{k\sqrt{e\delta}}{m-k}$. Thus the summation could be bounded by the integral of $f(x)$ on $(\frac{k\sqrt{e\delta}}{m-k},\infty)$
$\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\right]\Delta_{A_{t}}\right]$ $\leq 1540\sum_{i\notin S^{*}}\left(\int_{\frac{k\sqrt{e\delta}}{m-k}}^{\infty}f(x)dx+\frac{\ln ek^{2}}{\frac{k\sqrt{e\delta}}{m-k}}\right)$
$\leq 1540\sum_{i\notin S^{*}}\left(\frac{\ln\frac{\delta}{e^{2}(m-k)^{2}}-2\ln x}{x}\Big|_{\frac{k\sqrt{e\delta}}{m-k}}^{\infty}+\ln ek^{2}\sqrt{\frac{T}{e\log k\log(m-k)}}\right)$
$=1540\sum_{i\notin S^{*}}\left(\ln e^{4}k^{4}\sqrt{\frac{T}{e\log k\log(m-k)}}\right)$
$=O\left((m-k)\sqrt{\frac{T\log k}{\log(m-k)}}\right).$
Applying this lemma, we have
$\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$ $=\sum_{\ell=0}^{\lfloor\log(m-k)\rfloor}\sum_{t=1}^{T}\mathbf{E}\left[\left[A_{t}\in\mathcal{A}_{\ell}\cap\mathcal{A}_{B}\right]\cdot\Delta_{A_{t}}\right]$
$=O\left((m-k)\sqrt{\log k\log(m-k)T}\right).$
Bring this back to Equation 7, we could obtain that the regret of Algorithm 1 when $k>\frac{m}{2}$ is $O\left((m-k)\sqrt{\log k\log(m-k)T}\right)$.
@@
$=\begin{cases}O\left(\frac{\log k}{p^{*}}\sqrt{kmT}\right)&,k\leq\frac{m}{2}\\ O\left(\frac{m-k}{p^{*}}\sqrt{\log k\log(m-k)T}\right)&,k>\frac{m}{2}\end{cases}$ (9)
where Equation 9 applies Theorem 1 to this case.
For the conjunctive case, $r(A_{t},\bm{\mu})=\prod_{i\in A_{t}}\mu_{i}$. The rest analysis follows the same pattern as the disjunctive case.
@@ B.1 The CUCB algorithm
In Algorithm 3, if all $\omega_{i}(t)$ are less than $\left(\frac{1}{k}-\frac{\gamma}{m}\right)\sum_{i=1}^{m}\frac{\omega_{i}(t)}{1-\gamma}$, then every $p_{i}(t)$ calculated at Line 9 of 9 is less than $1$ and $S_{0}(t)=\emptyset$. Otherwise it chooses a appropriate threshold $\alpha_{t}$ that the temporal weight $\omega_{i}^{\prime}(t)$ is set to $\alpha_{t}$ for $i\in S_{0}(t)$ and $\omega_{i}(t)$ for $i\notin S_{0}(t)$. Thus $p_{i}(t)=1$ for $i\in S_{0}(t)$, which means that $S_{0}(t)\subseteq S(t)$. And $\omega_{i}(t)$ is not updated for $i\in S_{0}(t)$ since it is large to some extent.
@@ C.1 Comparison with baselines and beyond
As described in the main text, in the synthetic setting, we simulate a combinatorial bandit problem with $m=30$ base arms and cardinality $k=10$ and $k=20$. The initial value for each arm is independently drawn from uniform distributions $\mathtt{U}(0,0.1)$ and $\mathtt{U}(0.3,0.4)$. Each algorithm is run for $100000$ rounds, and all results are averaged over $10$ independent runs.
@@
To match the synthetic regimes used in our experiments, the raw affinity scores are min-max normalized and linearly rescaled into two ranges: a low-probability regime $[0,0.1]$ and a high-probability regime $[0.3,0.4]$. Specifically, let $\min$ and $\max$ denote the minimum and maximum dot products across all user–item pairs. The scores are transformed as follows:
@@
$\text{score}_{\text{high}}(u,a)$ $=0.3+0.1\cdot\frac{\mathbf{u}^{\top}\mathbf{a}-\min}{\max-\min}.$ (12)
The resulting values are interpreted as the initial click probabilities for base arms. For each regime, we independently sample 30 such values as base arms using a fixed random seed (2026) to ensure reproducibility.
@@
Table 2: Semi-bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.3,0.4]$, evaluated on both synthetic and Yelp datasets and fixed $m=30$.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Range $[0,0.1]$ Range $[0.3,0.4]$
Synthetic Yelp Synthetic Yelp
$k=10$ $k=20$ $k=10$ $k=20$ $k=10$ $k=20$ $k=10$ $k=20$
[…]
CUCB 4000.132 2348.979 4407.244 1916.620 3945.550 2361.077 4050.910 2214.263
EXP3.M 3523.649 1535.670 3094.234 1154.182 4009.885 2800.222 3555.937 2621.060
HYBRID 5858.858 3233.744 5197.661 2561.223 4706.725 2538.204 4067.236 2411.709
[…]
CUCB 15.155 27.816 11.422 21.108 10.724 20.391 11.739 23.882
EXP3.M 104.291 92.774 79.239 78.895 79.592 87.363 95.136 110.756
HYBRID 121.126 111.946 108.540 105.515 95.413 76.987 143.529 163.490
C.2 Ablation studies and beyond
  • Varying $k$: $k\in\{5,10,15,20,25\}$, with $m=30$ fixed.
  • Varying $m$: $m\in\{20,25,30,35,40\}$, with $k=15$ fixed.
On the synthetic dataset ($\mathtt{U}(0,0.1)$), CMOSS consistently achieves the lowest cumulative regret across all configurations. When varying $k$, CMOSS outperforms CUCB, EXP3.M, and HYBRID by up to 52.7%, 55.7%, and 67.4%, respectively. When varying $m$, the gains become even more pronounced—CMOSS reduces regret by as much as 53.3% over CUCB, 40.6% over EXP3.M, and 68.7% over HYBRID. In terms of runtime, CMOSS remains close to CUCB while being substantially faster than EXP3.M and HYBRID across all settings. Moreover, CMOSS exhibits stable runtime as $m$ increases, demonstrating strong scalability. On the Yelp dataset (rescaled to $[0,0.1]$), we observe consistent trends. When varying $k$, CMOSS achieves regret reductions of up to 51.0% over CUCB, 39.4% over EXP3.M, and 70.0% over HYBRID. When varying $m$, CMOSS again demonstrates clear advantages—outperforming CUCB by up to 46.9%, 23.1%, and 61.5%, respectively. Runtime comparisons further confirm CMOSS’s efficiency: it remains close to CUCB while being substantially faster than both EXP3.M and HYBRID. Additionally, CMOSS maintains stable runtime as either $k$ or $m$ increases, showcasing its practical scalability on real-world data. Complete numerical results are reported in Table 3 and Table 4.
Table 3: Ablation study of semi-bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on synthetic dataset. Results correspond to subplots (3)(4)(7)(8) in Figure 1.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Synthetic) Varying $m$ (Synthetic)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
[…]
CUCB 4429.727 3446.924 3184.634 2578.753 1131.007 1296.234 2182.487 2791.642 4208.573 5586.506
EXP3.M 4736.340 3271.894 2162.514 1462.691 659.203 836.489 1500.403 2327.747 3305.964 4369.808
HYBRID 5596.137 5123.518 4676.580 3456.369 1517.378 1918.722 3376.063 4285.878 6283.876 8184.837
[…]
CUCB 6.574 11.654 16.397 21.287 26.991 16.218 21.619 25.239 18.210 69.932
EXP3.M 76.208 120.789 107.957 116.241 106.603 61.540 96.724 124.403 129.031 447.195
HYBRID 125.178 135.696 108.276 91.548 88.490 63.963 96.495 147.526 172.044 245.882
Table 4: Ablation study of semi-bandit feedback: varying $k$ (with $m=30$ fixed) and varying $m$ (with $k=15$ fixed) on Yelp dataset.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Varying $k$ (Yelp) Varying $m$ (Yelp)
$k=5$ $k=10$ $k=15$ $k=20$ $k=25$ $m=20$ $m=25$ $m=30$ $m=35$ $m=40$
[…]
CUCB 4020.879 4208.577 3074.333 2175.718 1191.837 1162.516 2450.990 3299.031 3559.703 4558.815
EXP3.M 4021.209 2988.275 2025.406 1216.121 663.273 658.334 1409.273 2088.690 2424.766 3387.021
HYBRID 4969.380 5078.731 3960.924 2890.771 1944.711 1548.021 3338.523 4347.092 4913.297 5471.507
[…]
CUCB 7.448 13.630 19.281 25.876 31.900 16.236 24.373 23.632 28.706 71.921
EXP3.M 91.734 106.868 127.099 166.739 86.960 69.510 95.093 132.767 109.091 435.470
HYBRID 119.877 125.436 104.920 118.024 155.868 78.343 105.469 156.025 167.520 379.948
C.3 Parameter selection for EXP3.M
The performance of EXP3.M in semi-bandits is highly sensitive to the choice of its mixing coefficient $\gamma$, which controls the balance between exploration and exploitation. To determine an appropriate setting for our environment, we conduct an ablation over $\gamma\in\{0.001,0.01,0.1\}$ with cardinality $k\in\{10,20\}$. Each configuration is run for a horizon of $T=100000$ and averaged over 10 independent repetitions.
As shown in Table 5 and Figure 3, $\gamma=0.01$ consistently achieves the lowest cumulative regret across both initialization regimes. Under $\mathtt{U}(0,0.1)$, it reduces regret by $83.9\%$ and $80.3\%$ relative to $\gamma=0.001$ and $\gamma=0.1$ for $k=10$, and by $84.1\%$ and $44.8\%$ for $k=20$. A similar pattern holds under $\mathtt{U}(0.3,0.4)$: $\gamma=0.01$ improves over $\gamma=0.001$ and $\gamma=0.1$ by $75.6\%$ and $81.3\%$ for $k=10$, and by $67.3\%$ and $89.3\%$ for $k=20$. Based on these findings, we adopt $\gamma=0.01$ as the default parameter for all experiments involving EXP3.M, as it strikes the most favorable balance between exploration and exploitation in our stochastic environment.
Table 5: Ablation study of EXP3.M under different exploration parameters $\gamma$ on the synthetic dataset. Results are shown under two base-arm initial value distributions: (1) $\mathtt{U}(0,0.1)$; (2) $\mathtt{U}(0.3,0.4)$.
Metric Algorithm Range $[0,0.1]$ Range $[0.3,0.4]$
$k=10$ $k=20$ $k=10$ $k=20$
$\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$
Regret EXP3.M 21475.680 3457.137 17551.325 8570.484 1357.281 2459.607 17459.760 4255.882 22816.466 9514.021 3108.351 29068.987
Runtime (sec) EXP3.M 113.997 111.016 139.714 119.251 92.074 99.780 95.312 89.501 90.747 109.813 86.766 62.930
C.4 Cascading bandit
Furthermore, we simulate a disjunctive cascading bandit problem with $m=30$ base arms and cardinality $k=10$ and $k=20$. In each round, the algorithm recommends a list of $k$ items, and the user examines them sequentially—either in descending or ascending order of attraction probability—clicking on the first attractive item (if any) and ignoring the remaining ones. If none of the items is deemed attractive, the round results in no feedback and is treated as a no-op. The experimental configurations for both datasets (including synthetic and Yelp) and ablation studies in this section follow the same setup as detailed in Section 4 and Appendix C. Particularly, all experiments presented hereafter adopt this disjunctive feedback model. Visual comparisons across all experiments are summarized in Figure 4 (descending) and Figure 5 (ascending).
In addition, we do not include the configuration with $k=20,m=30$ under the $[0.3,0.4]$ regime. This is because, under the cascading setting, the regret gap becomes extremely small in this regime: with such a large slate size relative to the attraction range, the user almost always clicks on at least one item with high probability, making the optimal expected reward very close to 1 (as implied by the Bernoulli click model). As a result, different algorithms exhibit nearly indistinguishable performance, and the comparison becomes uninformative.
Descending bandit
[…]
Figure 4: Comparison of CMOSS (blue) with CUCB algorithms under cascading (descending) feedback. Subplots (1)(2)(5)(6) show cumulative regret with fixed $k=10$, $m=30$; (3)(4) show ablation studies varying $k$ (fixed $m=30$), while (7)(8) varying $m$ (fixed $k=15$). Subplots (1)(2)(3)(4)(7)(8) use synthetic dataset; (5)(6) use Yelp dataset. Initial means of base arms fall within the range $[0,0.1]$, except for (2)(6), which use the range $[0.3,0.4]$.
Table 6: Descending bandit feedback: cumulative regret and runtime (in seconds) of algorithms under two initial base arm value ranges: $[0,0.1]$ and $[0.3,0.4]$, evaluated on both synthetic and Yelp datasets and fixed $k=10,m=30$. Results correspond to subplots (1)(2)(5)(6) in Figure 4.
$\star$ indicates our proposed method (CMOSS).
Metric Algorithm Range $[0,0.1]$ Range $[0.3,0.4]$